OpenStudy (anonymous):

I don't get this,??? Suppose you select a card at random from a standard deck of cards 60 times, and 12 of those selections are hearts. How does the experimental probability compare to the theoretical probability?

Directrix (directrix):

How many heart cards are in a standard deck of cards?

OpenStudy (anonymous):

13

Directrix (directrix):

There are 52 cards in the deck. The 13 hearts are 13/52 = 1/4 the deck. Theoretically, if you draw 1 card from the deck, the theoretical probability of getting a heart is this: 13/52 or 1/4. Does that make sense?

Directrix (directrix):

@QueenTforlife Does that part make sense? We still have the experimental part to do.

OpenStudy (anonymous):

yeah

Directrix (directrix):

In the experiment, a card was drawn 60 times and 12 of those were hearts. The theoretical expectation is that 1/4 of the 60 draws would be hearts. What is 1/4 of 60 = @QueenTforlife

OpenStudy (anonymous):

15

Directrix (directrix):

According to the theoretical probability, you would expect 15 of the 60 cards to be hearts. But, in the experiment, 12 cards of the 60 cards were hearts. The experimental probability is 12/60 = 1/5

Directrix (directrix):

Theoretical Probability: 1/4 Experimental Probability: 1/5

Directrix (directrix):

How does the experimental probability compare to the theoretical probability?

OpenStudy (anonymous):

The theoretical probability is more than the experimental probability.

Directrix (directrix):

Yes. I'm thinking that you are also supposed to give the two probabilities as part of your answer.